Circular No. 5665

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

FG SAGITTAE D. Pollacco, Department of Physics and Astronomy, University of St.

Andrews, communicates: “CCD imaging by S. A. Bell, R. W. Hilditch, and D. L. Pollacco using the 1.0-m telescope at St. Andrews University Observatory shows that FG Sge has been slowly recovering from its recent unexpected deep minimum (IAUC 5604). Observations on twelve nights since Oct. 19 show an apparently linear increase of 0.015 mag/day (V) relative to five comparison stars in the field (11′ × 17′). The companion field star has not varied over this period by more than 0.015 mag and is therefore probably stable. These results appear to be consistent with an ejection and subsequent eclipse of material (cf. IAUC 5632) causing the abrupt minimum. Multicolor (BVRI) observations are available but await absolute calibration for the field. Observations are continuing and we encourage spectroscopic observations.”

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence,

communicate: “CCD photometry of V1333 Aql has been obtained at Haute-Provence since it went into quiescence around mid-June (IAUC 5551). Observations were made with the 1.2-m telescope on July 5-7 and 24-27, Aug. 4-8 and 18-24, and Sept. 1-3, 17-20, and 29-30. The source has remained in the quiescent state (V = 19) throughout this entire interval. However, in CCD frames taken on Nov. 27, the object appears bright (V = 16.7), at least a magnitude brighter than the average values (V = 17-18) seen during the last outbursts in 1989, 1990, 1991, and 1992 (see Chevalier and Ilovaisky 1991, A.Ap. 251, L11). This confirms the start of a new optical activity cycle coincident with the x-ray turn-on reported in IAUC 5664. Previous cycles have lasted ∼ 150 days and recurred every 300 days. This one has started roughly on schedule.”

NOVA CYGNI 1992 D. Hanzl reports his additional UBV photometry of N Cyg 1992 (cf.

IAUC 5623), obtained at the Nicholas Copernicus Observatory and Planetarium in Brno with a 0.4-m reflector (comparison star HD 195664): Oct. 14.745 UT, V = 10.14 ± 0.03, B–V = –0.60 ± 0.03, U–B = –1.19 ± 0.04; Nov. 7.724, 10.52 ± 0.04, –0.63 ± 0.03, –1.58 ± 0.02; Nov. 20.797, 10.62 ± 0.11, –0.47 ± 0.09, –1.74 ± 0.07.

1992 December 2 Daniel W. E. Green